OpenStudy (anonymous):

Which unit rate is the lowest price per ounce? Choice A: 25 ounces of potato chips for $3.49 Choice B: 45 ounces of potato chips for $4.79

OpenStudy (anonymous):

@phi

OpenStudy (phi):

price per ounce translate that to dollars per ounce or \[ \frac{dollars}{ounce} \] so I would do the division for both cases and see which gives the lowest dollars (meaning the cheapest) per ounce

OpenStudy (anonymous):

Choice A: 0.13 Choice B: 0.10

OpenStudy (anonymous):

right?

OpenStudy (anonymous):

@phi

OpenStudy (anonymous):

help!

OpenStudy (texaschic101):

make sure you round them correctly 3.49/25 = 0.1396 rounds to 0.14 (14 cents) 4.79/45 = 0.1064 rounds to 0.11 (11 cents) now which one is the cheapest ?

OpenStudy (anonymous):

oh, then the cheapest is choice B, right?

OpenStudy (texaschic101):

correct...it is the cheapest per ounce

OpenStudy (anonymous):

i have one more question. can you help?

OpenStudy (texaschic101):

I can try...can't make no promises...lol